OpenStudy (anonymous):

Sec^2x-1=0. Solve trigonometric equation

OpenStudy (mathmale):

Please take just a little extra time and write the trig function properly:\[\sec^2 x\] or 

OpenStudy (mathmale):

\[\sec^2x\]

OpenStudy (mathmale):

You have \[\sec^2x-1=0\] and are asked to find values of the angle, x, that satisfy this equation. Have you considered what you might do next?

OpenStudy (anonymous):

Add 1

OpenStudy (tkhunny):

\(\sec^{2}(x)\) \(\left(\sec(x)\right)^{2}\) Several options that exude clarity.

OpenStudy (mathmale):

Supposing we were to add 1 to both sides of this equation. What would happen, and would this result be any easier to solve? What would the next step be?

OpenStudy (anonymous):

Sec^2x=1. Then you square both sides??

OpenStudy (anonymous):

Directions say solve each equation over 0,2pi with the square root method. All answers must be exact in terms of pi.

OpenStudy (tkhunny):

You should probably consider factoring the Difference of Squares. Remember this? \(x^{2} - 1 = 0 \implies (x+1)(x-1) = 0\)

OpenStudy (anonymous):

Ohhhh ok

OpenStudy (anonymous):

But it says solve each equation with square root method

OpenStudy (mathmale):

The equation we're discussing is x^2-1=0. This is equivalent to x^2=1. Take the sqrt of both sides, remembering to write in a "plus or minus" sign on one side.